A spotlight S rotates in a horizontal plane with a constant angular velocity of 0.1 rad/s. The spot of light P moves along the wall at a distance of 3m. The velocity of the spot P when theta = 45 degrees is how much?

Grade:

3 Answers

Siddharth Mittra
12 years ago

Dear Sir,

I cant see any answer to my question. The place is blank. Please check and confirm.

Best rgds

Siddharth

promise
10 years ago
as w=d(theta)/dt=0.1 and x=3tan(theta) differentiate x wrt theta and now put theta=45degree multiply dx/d(theta) to w i.e.d(thta)/dt, you will get dx/dt=velocity ans to ur ques is 0.6m/s
Infinity Minds
5 years ago
Linear Velocity of rotating object ν=ωr/sinθω= angular velocity=0.1 rad/sr-distance of object from centre.angle=45°r=d/cosθ [ where d is the distance from the spot light to wall] =3/cos45°Velocity= ν=ωr/sinθ=[0.1x3/cos45°]/sin45°=0.1x3/sin45 °cos45°= 0.3/[1√2 x 1 /2]=2x0.3=0.6m/s∴The velocity of the spot P is 0.6 m/sRead more on Brainly.in - https://brainly.in/question/884832#readmore